OpenStudy (anonymous):

I need two coordinates for 5x+6y=30

jimthompson5910 (jim_thompson5910):

when x = 0, what is y?

jimthompson5910 (jim_thompson5910):

5x+6y=30 5*0+6y=30 ... plug in x = 0, ie replace x with 0 0 + 6y = 30 y = ??